### v8

# Population of tetraneutron continuum in reactions of 8He on deuterium

Abstract
Search for the population of the low-energy continuum of a tetraneutron system was performed for reactions of the 8He beam on a deuterium target. These studies are based on the data [I.A. Muzalevskii et al., Phys. Rev. C 103, 044313 (2021)], previously used for the studies of 7H and 6H in the H and H reactions. Evidence for a hump in the 4n continuum at and MeV was observed in the 2H(8He,6Li)4n and 2H(8He,3He)7HH+4n reactions, respectively. The observed statistics is quite low (6 events and up to 40 events) corresponding to very low cross sections of few microbarns or tens of microbarns. The background conditions for the 2H(8He,6Li)4n reaction are shown to be good, favoring the physical nature of the observed events. The 2H(8He,3He)7HH+4n process transforms to the 2H(8He,6Li reaction in the limit of the highest 7H decay energies. The population of the low-energy region in the 4n spectrum is found to be perfectly correlated with the population of the lowest 6Li state in the 3He+3H continuum with MeV. Theoretical calculations of 8He in a five-body + and of 4n in a four-body hyperspherical models are presented. The 8He wave function is shown to contain strong specific correlations, which may give rise to very low-energy structures in tetraneutron continuum in extreme-peripheral reaction scenarios.
I Introduction
The search for multineutron systems is old, but still unsettled problem of the low-energy nuclear physics. The first ideas about the possible nuclear-stability of multineutron systems were expressed in Refs. [1, 2] and since that time reexamined on various occasions. A detailed account of the search for bound multineutron, both experimental and theoretical, can be found in the recent review [3]. In short, either none or only marginal experimental evidence, which was never confirmed later, was obtained for the bound tetraneutron. However, such attempts continue, and there is a very recent example [4]. The situation has changed with the recent studies of the low-energy 4n continuum population in reactions with 8He, where four neutrons can be found in a spatially-separated neutron-halo configuration. In the 4He(8He,) reaction [5] 4 events were observed in the low center-of-mass (c.m.) 4n energy range MeV with the tiny cross section of nb. Even more recently, a statistically convincing peak (reported as “resonance-like structure” comprising around 44 events) was observed in a 1H(8He,) experiment [6] at MeV with MeV. The reported cross section in this study is also small and estimated to be sub-microbarn. It should be noted that the both observations [5, 6] belong to the very specific situation of extreme backward (backward/forward in the case of [5]) kinematics of quasi-free scattering.
The issue of bound “multineutron nuclei” was scrutinized in the modern theoretical approaches [7, 8, 9] with the same result: for fixed two-body potentials, radical modifications of the three-body potential are required to get bound 4n. Such modifications are considered unacceptable as they are inconsistent with well-known nuclear structure of all other nuclides. In the theoretical studies of the low-energy 4n continuum various indications were found for important low-energy effect of the 4n final-state interaction (FSI). These include: specific trajectories of the -matrix poles, enhanced “time delay”, energy extrapolations of multi-neutron states confined in traps, etc., see e.g. [10, 11, 12, 13, 14, 15, 16]. There have been, however, two works that have predicted a 4n resonant state [17, 18], strongly contradicting the other studies. It has been mentioned above that the illuminating summary of those studies can be found in the review [3].
An approach to 4n system tractable in terms of observables (namely, the reactions with neutron-rich halos) was presented in Ref. [19]. In this work the 4n continuum is populated from the “prearranged” 4 neutron configurations in the “atmosphere” of a halo nucleus 8He. The approach predicts population at low energies, but this population is a result of a specific (very peripheral) initial state structure (ISS) sampling, modified by FSI and should not be considered a stand-alone resonance. The formation of such a low-energy continuum response is partly analogous to the mechanism of the “soft dipole mode” formation in halo nuclei: (i) relatively very weak FSI effects, but nevertheless, strong low-energy concentration of the strength function due to (ii) the peripheral character of the halo wave functions (WF) and (iii) the further periphery enhancement by electromagnetic transition operators [20, 21, 22]. Principal message of [19] is that both halo character of 8He WF and 4n FSI are not yet sufficient to produce low-energy response in the 4n continuum, and additional assumption about peripheral nature of the reaction mechanism is required.
In Ref. [23] the data of [6] were analyzed under assumptions similar to Ref. [19]. However, conclusion was made in [23] that the 8He-induced source can directly explain the data of [6] without any peripheral assumptions on reaction mechanism. So, there is a qualitative contradiction between the results of [19] and [23] based on more or less the same physical assumptions.
In this work we demonstrate that an evidence for the low-energy structures, analogous to the observations of [5, 6], can be found in the other reactions with the 8He beam. As for the theoretical discussion, we revisit the 4n population from 8He in simple reaction approach of Ref. [19] (sudden removal approximation) introducing, however, important technical improvements. We generally confirm the results of [19], but provide an additional insight into the reaction mechanism, connected with specific correlations in the 8He WF. Possible reasons for disagreement with the results of [23] are analyzed.
The system of units is used in this work.
II Experiment
Experiments with the 8He beam impinging on the deuterium target was performed at the ACCULINNA-2 facility (FLNR, JINR) having the search for the 7H nuclide as a goal [24, 25]. The experiment appeared to be much “richer” than initially intended, providing interesting results also on 6H [26] and other auxiliary channels like 8Li, 9Li [27, 28] and 5H, 5He, 7He [29]. Here we present the results obtained for the 2H(8He,6Li) and 2H(8He,3He)7HH+4n reactions, previously omitted. Since the experiment has been already well presented in literature, we only briefly describe here the most relevant details.
The 8He beam was produced at 26 A MeV with mean intensity around pps. The beam was focused on the cryogenic gaseous deuterium target, with a temperature of 27 K, equipped with the thin stainless-steel and mylar windows. The target thickness was . For the 2H(8He,6Li)4n reaction the setup allowed us to measure the 6Li recoil in coincidence with a neutron from the produced unbound 4n system.
The beam particles were identified by the two plastic scintillators, allowing to deduce the energy of the projectile from its measured time-of-flight (ToF). The projectile trajectories were reconstructed by the two pairs of multi-wire proportional chambers. A special run with empty target cell was performed to estimate the background conditions, which was of the total 8He beam time long.
The recoil 6Li nuclei, appearing in the 2H(8He,6Li) reaction hit the array of four identical -- telescopes. The telescope array was located 179 mm downstream the target. Each telescope consisted of three layers of silicon strip detectors (SSDs). The 20-m-thick SSD with a sensitive area of mm2 was divided into 16 strips, the second and the third layers consisted of the two identical 1 mm-thick SSDs ( mm2 with 16 strips). The 6Li recoils emitted from the deuterium gas target in the 2H(8He,6Li) reaction were detected by this telescope array in a range in the lab system. The telescopes allowed to identify 6Li with clear separation from the other registered lithium isotopes, see Fig. 1 (a). The central telescope was installed at the beam line at the distance of 323 mm behind the target. It was intended to detect tritons emitted with high energies at the angles smaller than in the lab system. The telescope consisted of one 1.5-mm-thick double-sided SSD ( mm2, with 32 strips on each side) followed by a square array of 16 CsI(Tl) crystals. The crystals had a cross section of mm2 and thickness 50 mm each.
The group of four neutrons appears in “free flight” as a result of the -core removal from the 8He projectile. An important part of the 2H(8He,6Li)4n reaction analysis was the neutron identification and reconstruction. The neutron-wall setup [30] included 48 stilbene scintillator crystals placed on a m2 area located 2 m downstream the target at zero angle to the beam axis. The distance between the 50-mm thick and 80-mm diameter stilbene crystals was approximately 12 cm which resulted in about probability for neutrons to hit a stilbene detector. The stilbene array provided energy resolution and the cumulative single-neutron registration efficiency of . The probability of a neutron registration in coincidence with the 6Li recoil can be estimated as , taking into account that four neutrons are flying forward, towards the stilbene array. The -gamma separation was made by means of the Pulse Shape Discrimination (PSD) [30]. The PSD information, supplemented with ToF distribution, see Fig. 1 (b), leads to suggestion that some gamma-type signals correspond to neutron-like ToF. These gammas are presumably produced by neutrons interacting with the stilbene modules housings and should be taken as neutron events.
II.1 2H(8He,6Li)4n results
Tetraneutron spectrum was reconstructed from the recoil 6Li as a missing-mass (MM) component in the 2H(8He,6Li) reaction. For background reduction we considered 6Li coincidences with one of the neutrons coming from the 4n decay. The total number of 6Li- coincidences found in the recorded data was 136, see Fig. 2 (a). In this figure the neutron kinetic energy in the 4n c.m. frame is plotted opposite the reconstructed MM energy of the group. The fact that the majority of these events (108 events) are located inside the kinematically allowed region confirms a good channel identification for the data. The empty-target measurement Fig. 2 (a) shows that the background conditions were very “clean” for the 4n population in the 2H(8He,6Li) reaction: the empty-target events are very few and located mainly in the unphysical part of the kinematical plane.
The MM spectrum of 4n, Fig. 2 (b), was reconstructed from the measured 6Li recoil energy and emission angle, taking the 108 events located within the “kinematical triangle” . The obtained spectrum shows a group of 6 events around MeV. The width of this group is reasonably consistent with the value of energy resolution MeV FWHM in this region of 4n MM spectrum. The green dotted curve in Fig. 2 (b) shows the Monte Carlo (MC) simulated empty-target contribution. The statistics of the low-energy group of events is quite low, but the energy resolution is sufficient and the background conditions are shown in Fig. 2 to be very good. It can be evaluated that less than 1 random event is expected in the energy range MeV. Then the Poisson probability to get 6 random coincidences in that energy region is .
Obviously, the observed low-energy hump in Fig. 2 (b) can not be described just as a contribution of the 4-body phase volume or Fourier transform of the 8He source, see also Fig. 8 (d) and the related discussion in Sec. VI. The phase volume for uncorrelated decay of -body system, which can be regarded here as a “minimum admissible model”, is
and the generic 4-body phase volume behavior is . However, Eq. (1) should be modified by the minimum number of excitation quanta allowed for the continuum WF as
For the 4n system the Pauli principle requires two additional excitation quanta for the lowest-energy configuration. This corresponds to lowest-energy configuration in the shell-model terms or to in terms of the hyperspherical harmonics (HH) model principal quantum number (hypermomentum), see Section IV. This leads to a modified 4-body phase-volume behavior shown in Fig. 2 (b).
It is notable that the mean energy MeV of the low-energy group of events in Fig. 2 (b) is consistent (within experimental energy resolution) with the 4n low-energy peak value MeV reported in Ref. [6].
The 2H(8He,6Li)4n reaction has already been used before for tetraneutron search, but the results were not published in details, only briefly in conference proceedings [31, 32]. In Ref. [31], a resonance-like structure (with quite a low statistics) was reported at MeV. In contrast, for the next analogous experimental run [32] the 4n spectrum with MeV was found to be dominated by the carbon background due to the use of the CD2 target. As compared to studies of [31, 32] our experiment has no carbon background and additional strong background elimination, connected with kinematics of Fig. 2 (a), is possible.
Another background issue to be discussed is possibility that the 2H(8He,6Li)4n reaction results could be contaminated by the 2H(8He,n reaction contribution, where is the isobaric-analogue state of the 6He(g.s.). Since the state decays only electromagnetically, we cannot distinguish the latter reaction contribution in the current experiment.
Fortunately, there is a strong argument why the population of should be very small. It is well established that reactions at typical energies of our experiment (26 ) and at higher energies are dominated by single-step direct-reaction mechanism. Within this paradigm the 4n subsystem is “spectator” (inert during the reaction time) and transfer is induced by the T-matrix between the 4He “participant” and the deuterium of the target:
where is sum of all possible nuclear interactions in the channel of interest. However, it is well known that the state can be obtained in 2H capture on -particle only by electromagnetic M1 transition or by parity-non-conservation neutral-current weak interaction:
where the parity-violating nuclear interaction is known to be extremely small. Indeed, the cross section for the capture reaction 4He(,Li has been measured in a broad energy range by detecting gammas or recoil 6Li ions [33]. No 3.56 MeV resonance population has been observed and the best limit on the parity-forbidden width of this state was obtained as tiny eV in Ref. [34]. Thus, the 2H(8He,n reaction may be connected only with two-step and more complicated reaction scenarios, which are expected to be suppressed.
It should be also understood that actual population will not dismiss the low-energy 4n issue, just make its interpretation more complicated. Assume extreme situation that only was actually populated. This means that we miscalculate the missing mass of 4n, which should be actually 3.56 MeV smaller according to energy-conservation. So, the low-energy 4n events in Fig. 2, which have energies between 2.5 and 5 MeV after recalculation will have energies from to 1.5 MeV with centroid at around the 4n threshold. Existence of such a “severe low-energy” 4n correlation, is even less realistic and complicated to interpret theoretically.
II.2 2H(8He,3He)7HH+ results
In Ref. [25] the excitation spectrum of 7H was populated in the 2H(8He,3He)7H reaction up to MeV above the 5-body 3H+ decay threshold. An evidence was reported for resonant states at and 5.5(3) MeV and also some indications for such states at 7.5(3) and 11.0(3) MeV. These structures were observed at small c.m. reaction angles . The 7H excitations with MeV observed in a broad angular range form quite a featureless hump. It is possible to search for the low-energy-correlated 4n emission off these sufficiently highly-excited configurations of 7H.
The data of Ref. [25] contain coincidence information on the 3H decay product of 7H, which allows one to infer the MM of the 4n subsystem. These data in terms of the 4n emission are shown in Fig. 3 (a). The 7H events are mainly concentrated near the diagonal , which means that the “internal” energy of the tetraneutron tends to be large, while the triton gets only a small fraction of the total 7H decay energy. This situation is close to the phase-volume distribution of the 5-body 3H+ decay for the given 7H decay energy as
in terms of 4n energy-distribution parameter
The standard value for the 5-body phase volume is . Eq. (3) is the double-differential form of the Eq. (2), discussed in the previous Section, so it can be found that at least value is requested for by the Pauli principle for 4n. This idea is confirmed by the calculations of the 5-body decay of 7H in [35]. The distributions at given 7H energies are also quite “compressed” towards small 3H energies due to the experimental bias, see Fig. 13 of [25], so for the distribution of Eq. (3) turns out to be completely justified in our situation.
To observe a possible low-energy 4n correlation, we need to find kinematic region, where the main “phase-volume” component of the decay distribution Eq. (3) is well suppressed. This means , and, therefore, for studies of the 4n excitation region MeV we need to consider the 7H decay energy region MeV. Under this selection condition, the hump with MeV can be found in the 4n MM spectrum, see Fig. 3 (b). The MC simulated phase volume only weakly “penetrates” in the MeV region. The background of about events/MeV is estimated from the available empty target data, see green circles in Fig. 3 (b). Thus a significant part of this low-energy 4n data may be connected (but not necessarily is connected, as we see below) with a random background.
The 7H data can be converted to presentation of system reconstructed as 3He+3H continuum, see Fig. 4 (a). Fortunately, it appeared that the low-energy 4n events in Fig. 3 (b) are practically perfectly correlated with the population of the lowest state in the 3He+3H continuum, with MeV and MeV [36], see Fig. 4 (b). This means that we are actually dealing here not with the decay, but rather with the 2H(8He,6Lin reaction. Moreover, the same correlation is perfectly true not only for events in the MeV range, but also for the MeV range. In this range most of the events should be either unphysical or connected with insufficient energy resolution, as the population of the MeV range is predicted to be negligible in available theoretical scenarios, see Fig. 7 or Ref. [23]. The MC estimated energy resolution for the 4n MM spectrum is around MeV FWHM. However, this resolution has a broad “nongaussian” component, which may result in some number of events with MeV. The coincidence with the definite state in 6Li practically guarantee the physical nature of such event, so we recognize that the absolute majority of events (26 out of 28) found in the range MeV in Fig. 3 (b) as really belonging to the 4n MM spectrum.
Strong correlation found in Fig. 4 (b) inspired us to “reverse” the logic and construct the 4n spectrum gated by the lowest states in the He+3H continuum, see Fig. 4 (c). The profile of the low-energy MeV 4n spectrum is practically insensitive to the gate energy selection for MeV. For higher energies the 4n spectrum expectedly becomes phase-volume-like. The peak energy can be found as (sys.) MeV. That peak position is consistent with (sys.) MeV of Ref. [6] within the experimental energy resolution. Statistically our data is analogous to data [6]: 44 events in [6] vs. 26 and 40 events in the and MeV gates, respectively.
There is a strong sensitivity of the higher-energy part of the obtained 4n MM spectra to the energy of the populated 6Li states. For example, it can be seen in Fig. 4 (c) that 4n spectra in the gates and MeV show very different populations of 4 and 25 events in the MeV range, which seems to be beyond statistical uncertainty. This may be indication that mechanisms for 4n population within these gates are somewhat different: the states of the 6Li recoil are broad and overlapping in this energy range [36], and the higher-energy gate could be “contaminated” by contributions of the higher excitations of 6Li.
III On tetraneutron population in different reactions
The low-energy population of 4n in Ref. [6] and in the processes considered in this work all have the 8He nuclide as a starting point. At first glance, this is the only similarity, and the reaction mechanisms should be too different. Let’s demonstrate that the differences are actually not so striking.
Table 1 shows the momenta and velocities in the recoil-product channel for the reaction of Ref. [6] and the reactions considered in this work. Of course, in the reaction of [6] the “recoil” particles are leaving the FSI region faster, than in the case of our Li) reaction, but in the timescale this “faster” is only a factor of 2.
| Reaction | Recoil | (MeV) | |
| 1H(8He,)4n, Ref. [6] | 5Li | 750 | 0.36 |
| 2H(8He,6Li)4n | 6Li | 426 | 0.19 |
| 2H(8He,3He)7H | 3He | 311 | 0.16 |
| 7HH+4n | 3H |
The 2H(8He,3He)7HH+4n process formally has the 7HH+4n decay process as its second step. In fact, from Table 1 it is clear that for the relatively highly-excited states of 7H, which we consider, the velocity of the 3H recoil leaving 4n in the “decay process” is actually comparable with the velocity, at which the 3He recoil had left the FSI region at the first [the 2H(8He,3He)7H reaction] step of this process. This kinematical region is actually a transition region from 2H(8He,3He)7H reaction to 2H(8He,n reaction, where highly excited states of 6Li (those located above the 3He-3H threshold) are populated. We find in Fig. 4 (c) that the most pronounced low-energy 4n hump, not “contaminated” by higher-energy phase-volume-like contributions, is obtained for the recoil, which is the lowest-energy resonant state in the 3He-3H continuum.
So, we see that the low-energy 4n continuum is confidently populated in the 2H(8He,n reactions with excitation energies MeV, see Fig. 2 (b), and MeV, see Fig. 4 (c). It is quite natural to expect that such a population should also take place at the intermediate energies, which correspond to + continuum states of 6Li. In this case we deal with reaction of quasifree scattering 2H(8He,)4n, which has never been considered seriously because of extremely “soft” nature of the 2H target. Unfortunately we were not able to extract this information from our existing data due to technical limitations.
Another characteristic, which is expectedly very important for the direct reactions, is momentum transfer to the 4n system, see, e.g. the Eq. (9). For our reactions the reconstructed values are shown in Fig. 5. Quite different momentum transfers and MeV are realized for the low-energy 4n population in the 2H(8He,6Li) and 2H(8He,3He)7HH+4n reactions, respectively. This is found important for data interpretation, see Fig. 7 and related text. Unfortunately, we do not find information on the momentum transfers in Ref. [6], though this information is obviously accessible in these data. Only the restriction MeV can be roughly estimated from Fig. 2 of Ref. [6], while the most probable value seem to be around MeV.
IV Theoretical model
| Th. | ||||||
| Exp. |
In this work we performed studies of 8He in a five-body + and of 4n in a four-body hyperspherical harmonics models. A detailed description of these studies will be given elsewhere, while here we focus on the aspects related to 4n production from 8He, following the ideology of Ref. [19]. The HH Schrödinger equation (SE) for particles is used either without (bound states) or with right-hand-side inhomogenity (continuum states):
On the properly antisymmetrized HH basis the variational procedure reduces the SE with inhomogeneous term (4) to a set of coupled ordinary differential equations:
where are two-body potentials or , while are three-body potentials or . The “multiindex” is numbering the available HH basis states. The hypermomentum is the grand quantum number in the few-body configuration space and
For 8He and 4n its minimum value is due to the Pauli principle for four neutrons.
For the SE (5) without Coulomb interaction the boundary conditions at large are known analytically
where , are the modified Bessel functions of the second kind, are the Riccati-Bessel functions.
The “source functions” are terms of hyperspherical expansion of the inhomogeneous term :
The simplest reaction model which can be considered for 4n production is “sudden removal” of core from 8He [19, 23]. This procedure defines the source function in (4) as the Fourier transform of the overlap integral of the -cluster WF and the 8He WF over the radius-vector between the removed -cluster and the 4n center-of-mass:
Within the sudden removal approximation the momentum transferred to the 4n system is equal to the momentum of the removed -cluster. The “strength function” for population of the 4n continuum should be proportional to the outgoing flux of particles, expressed via the WFs with pure outgoing wave boundary conditions Eq. (7)
on a hypersphere of a large radius fm.
V Properties of the 8He
Following the cluster ++ model studies of 6He, 6Li, 6Be isobar [37, 38, 22] we use the Sack-Biedenharn-Breit (SBB) potential [39] in the - channel. We suppress the Pauli forbidden state by using the additional repulsive core in -wave potential, which provides practically the same scattering phases as the deep potential.
| Reaction | 8He-, | 8He-, | 8He-12C, |
| Th. | |||
| Exp. | [40] | [40] | [41] |
In the calculations we adopted the local charge-independent Argonne SSC AV14 potential [42]. For evaluations we also used simple -wave singlet potential Brown-Jackson (BJ) from [43]: , MeV, fm. For three nucleons with separations we use the three-nucleon potential [44]
in Eq. (6). A set of the parameters MeV, , MeV, yields the binding energies 8.41 (8.48), 7.74 (7.72), and 28.44 (28.30) MeV for 3H, 3He, and 4He ground states, respectively (the experimental values are shown in parentheses).
We also adjusted the 4He-- three-body potential of the same form as Eq. (11) to reproduce the ground state properties of 6He, which is well described in terms of the three-body 4He-- model [37, 38, 22]. For 8He g.s. the calculations were performed with , containing 15862 basis states which are reduced to 675 states in (5) by antisymmetrization. The 8He is slightly underbound (by approximately 115 KeV). This means some overbinding at higher , but this does not matter for the study of the tetraneutron: the geometry of the 8He g.s. WF is much more important here. The percentages of the major 8He WF components are in a good agreement with advanced ab initio Quantum Monte Carlo calculations of [45] and [46] . The geometric characteristics of 8He are compared with experimental data in the Table 2. The root mean square (rms) charge radius of 8He, has been determined for the first time as 1.93(3) fm [47]. However, using the new charge radius 1.67824(83) fm of 4He [48] as an anchor point for the isotope shift measurements, the 8He charge radius was reevaluated as 1.9559(7)(158) fm, where the first uncertainty is from new charge radius value and the second uncertainty from the electronic isotope shift measurements. Our result for 8He charge radius is in nice agreement with both experiments. The comparison with experimental rms matter radius in Table 2, seems to be not so favorable, but we need to recall that this value is actually extracted from the high-energy reaction data in a model-dependent way. Therefore, here the direct comparison of cross sections is more preferable. The results of Glauber-like calculations with 8He densities obtained in our model, shown in the Table 3, are in a very good agreement with experimental data, including the cross section on the 12C target [41], from which the matter radius of fm, shown in Table 2 was actually extracted. Thus, we can expect that our 8He WF, which near perfectly describes all known experimental data and reveals extremely important -4n correlations is a reliable starting point for the tetraneutron study.
VI The 8He WF as a source for tetraneutron population
The 8He WF correlation density vs. (collective radius of four neutrons, informative for the 4n studies) is shown in Fig. 6 (a). It looks overall quite featureless. In contrast, the projection of the 8He WF on the 4n configuration with , where strongest 4n FSI effect is expected, exhibits strong spatial correlations, see Fig. 6 (b). The correlation density vs. can be easier perceived basing on a simple relation between the rams radius of individual neutron in the 4n c.m. frame and the 4n rams hyperradius
Following the well-known naming of spatial correlations in the classical 6He halo nucleus — “dineutron” and “cigar” [37] — we characterize correlations in 8He source as “broom” (compact tetraneutron aside of -particle) and “cross” (neutrons evenly distributed around central-located -particle). The double-hump picture of correlations can be qualitatively understood as connected with configuration (in the shell model notations) in analogy with configuration providing strong spatial focusing (“Pauli focusing” [37]) in the 6He case. Fig. 6 (c) shows the Fourier transform of the source (b), illustrating the momentum content of this structure without 4n FSI. So far, the Pauli focusing for the 8He g.s. WF has been several times discussed in the qualitative models [49, 50, 35], but has never been demonstrated in the realistic calculations of the 8He g.s. WF.
It should be understood that the strongly correlated source shown in Fig. 6 (b) is related to about of the 8He g.s. WF. The other components of these WF projected on 4n system populate “excited” configurations of 4n, where we do not expect any strong FSI effect.
Figure 7 (a) shows the evolution of 4n strength functions for one selected MeV: no FSI reduced FSI full 4n FSI. The strength function peak for full 4n FSI is at about 14 MeV in a good agreement with analogous calculation in [19]. To form extreme low-energy peak in the strength function in the region MeV as indicated by data, a further qualitative assumption of the peripheral character of reaction is needed. To evaluate it, we cut out small hyperradii of the original source function by a profile function of the form
and further refer it as “peripheral source”. With fm and fm the rms size of the source is increased to fm and the desired energy position of the peak is obtained. It should be noted that the low-energy peak in the latter calculation is associated with “cross” component of the source, while the higher-energy peak at about MeV can be related to the “broom” correlation.
The systematics of the strength functions for peripheral sources for different momentum transfers is illustrated in Fig. 7 (b). One may see that even the “peripheral assumption” is not sufficient to form a strong and distinct low-energy peak in the 4n spectrum: the desired picture is realized only for certain intermediate range of transferred momenta MeV. This range well corresponds to situation of our experiment, see Fig. 5 and the related discussion in Section III.
The calculations for 4n are performed up to , which is insufficient for complete energy convergence. For that reason we perform simultaneously the calculations with advanced AV14 potential and simple central - potential BJ with much faster and “safer” convergence. The convergence curves are used for exponential extrapolation to infinite basis. For example, for calculations of Fig. 7 (a), the convergence point was MeV, while “underbinding” was MeV for AV14 and MeV for BJ potentials. For Fig. 9 the corresponding values are , and MeV. The additional control of the energy may be provided by phenomenological few-body potential with Fermi-function profile
The quite “soft” 4-body potential (13) with parameters fm and fm was used to get the peaks at energies pointed by the exponential extrapolation. “Safety” of the procedure in the sense of strength function shape was checked by extrapolations.
Qualitative comparison of theory and available data is provided in Fig. 8. Here we need to recall that strongly correlated picture of Fig. 7 is connected with of the 8He WF projecting on the state of 4n, while of this WF populate excited configurations of the 4n. The latter WF components are evaluated to provide strength function with maximum at MeV, see thick gray curve in Fig. 8 (d). As we have mentioned in Section III, the information on transferred momenta is not available in Ref. [6], only the rough upper limit estimate MeV. This means that, so far, we can not find out even on the qualitative level, whether the experimental results of this work should be in principle, the same as in [6] (and they just seem somewhat different only because of experimental resolutions and statistics), or there should be some significant physical difference between them due to difference in the reaction mechanisms connected with value.
It is also important to note that the low-energy peak connected with extreme peripheral source has very different profile compared to the case of the real resonant behavior connected with normal-size source. The “resonant” (blue dashed) curve in Fig. 8 (d) was generated by additional artificial binding introduced in Hamiltonian by phenomenological 4-body potential (13) with parameters fm and fm. The difference between “peripheral” black solid and “resonant” blue dashed curves in Fig. 8 (d) is too small to be seriously discussed for the available quality of the data. However, in principle, we have guidelines from theory how to distinguish experimentally the “peripheral” (caused by reaction mechanism) and “true resonant” tetraneutron continuum responses.
VII Theoretical discussion
In the recent paper [23] the data of [6] was qualitatively reproduced by emission off 8He-induced source without any peripheral assumptions. This is in a strong contradiction to the results of [19] and this work. Let’s try to track the origin of this contradiction. In Fig. 3 of Ref. [23] a calculation with reduced - interaction (scattering length fm) is shown. This calculation is reasonably close to “no FSI” situation and thus its origin could be to easier to “trace”. It is shown in Fig. 9 that a very close result is obtained in our calculations with simple analytical source
for the , component. The rms neutron radius fm of such a source is very consistent with the values found in Table 3 of Ref. [23]. The strength function peak shape is somewhat different from that in Fig. 3 of [23], which indicates more complicated correlated nature of the source function in Ref. [23], then we assume in Eq. (14). However, the peak value for the full-calculation 4n strength function in Fig. 9 well coincide with 2.5 MeV of [23], which means that we are reasonably “on a right track”.
So, in principle, there exist a class of source functions, which can produce extreme low-energy peak in the 4n spectrum for quite a compact ( fm) source. The only problem is that the “supersoft” source Eq. (14) has short-distance asymptotic . If we get back to the definition Eq. (8), this implies that 8He overlap with 4-body hypersphericals is singular as at zero, which is highly unlikely for any realistic 8He structure. The 4n sources deduced in our work has asymptotic at zero for ; they can be reasonably approximated by expression like (14), but only for . This is in full agreement with analytically deduced source functions of Ref. [19]. The source with is evidently much “harder” in the momentum space than the source of the same radius with . Therefore, an additional peripheral assumption is needed to get sizable low-energy effect in 4n spectrum with such a source.
So, we guess that the low-energy enhancement of the 4n spectrum in Ref. [23] is connected with “supersoft” source. The origin of the “supersoft” source in this work is probably due to the fact that 4n c.m. is assumed to coincide with 8He c.m. in Ref. [23] leading to several important inconsistencies:
(i) This is an implicit assumption of infinitely heavy -core in 8He and, of course, quite a poor approximation considering the actual mass ratio of and 4n.
(ii) This also means that within the cluster model the charge radius of 8He coincides with the charge radius of the -particle, which contradicts to the experimental data [47].
(iii) The “infinitely heavy core” assumption also leads to total absence of such an important characteristic as in the formalism of [23]. In contrast, the “geometric” effects of 4n motion relative to in 8He WF are found to be very pronounced in our calculations, see Fig. 6.
(iv) The rms size of the source fm, see e.g. Fig. 7 (a), presume significantly smaller fm distances in the 4n source, than the fm distances in the initial 8He WF, see Table 2. The “infinitely heavy core” assumption leads to equal distances in the 4n source and in the initial 8He WF.
So, we insist that peripheral assumption is necessary to describe the extreme low-energy peak in the spectrum of 4n. This point is also qualitatively supported by extreme small cross sections for the observed peculiarities in the low-energy 4n spectra. Typical cross section values for analogous direct reactions are some millibarns and even tens of millibarns, while we and Refs. [5, 6] observe some microbarns and even nanobarns. Such a cross section suppression looks quite natural if only the extreme periphery of the original 4n configuration of 8He is actually participating in the reactions of interest without destructive rescattering effects.
VIII Conclusion
Evidence for a hump in the continuum at and MeV was observed in the reactions 2H(8He,6Li)4n and 2H(8He,3He)7HH+4n, respectively. The 4n peak energies in our work and the “resonance-like structure” at MeV observed in Ref. [6] are consistent within experimental resolutions. We demonstrate that the reaction mechanism in both reported here reactions is actually best interpreted as transfer reaction to the ground state and to the MeV excited state of 6Li, respectively. The reaction mechanism also can be found reasonably consistent with that in the 1H(8He, 4He)4n reaction [6], so it could be the same type of phenomenon. The statistics of 6 events for the first our reaction is quite low, but events for the second one is comparable to statistics connected with the low-energy 4n resonance-like structure observed in [6].
Paper [23] is explaining the low-energy tetraneutron peak in Ref. [6] as solely an initial state (8He structure) effect in the presence of 4n FSI. This is in a strong contradiction to the results of [19] and the theoretical results of this work. As compared to [19] the calculations of this work (i) are performed with realistic and interactions, and (ii) the obtained realistic 8He source demonstrates the complicated correlated behavior that is very important for the 4n population. So far, the strong spatial correlations (Pauli focusing) for the 8He g.s. WF has been several times discussed in the qualitative models [49, 50, 35], but in our work they are for the first time explicitly demonstrated in the realistic calculations of 8He g.s. structure.
The two aspects mentioned above are not sufficient to fully explain the presence of a low-lying peak in the tetraneutron spectrum. To do this, it is also necessary to consider extremely peripheral reaction mechanism, in which the -core of 8He interacts with the target, but four neutrons remain “spectators” at distances, far exceeding the typical sizes of neutron orbitals in 8He. Using the appropriate profile function of the type Eq. (12), simulating such a geometry, it is possible to explain the presence of a low-lying peak in the tetraneutron spectrum. Since the appearance of the low-energy peak is not related to the tetraneutron per se, but also to the ISS (the source used) and the reaction mechanism, it should be expected that such a structure could be detected at somewhat different energies in different reactions. It is demonstrated that a likely reason of disagreement with the calculations [23] may be the incorrect 4n c.m. treatment in the 8He WF used in [23].
The existence of low-energy tetraneutron resonance would mean a radical revision of everything we know about neutron-rich nuclei and neutron matter. Our vision of the problem is that a solution can be found, which is much less radical being related to the 8He structure and reaction mechanisms. We got an important indication in our data that the 4n spectrum is strongly dependent on the particular state of the recoil 6Li system, which means, first of all, on the reaction mechanism. If our understanding of the low-energy structures in tetraneutron is true, then it is actually not disappointing at all not to find tetraneutron resonance: we get here an important cautionary lesson concerning low-energy observables in exotic nuclei and (for those interested) we arrive at a new exciting field of studies of extreme-peripheral phenomena in nuclear systems.
It is clear that the further studies of 8He induced reactions on the deuteron target are needed to bring to full reliability the evidences obtained in this work. However, the processes of the major interest are all in the cross sections range of microbarns. Therefore, such studies make sense only if the statistical limitations of the existing data are overcome by about an order of magnitude, which is a real challenge for available experimental techniques and facilities.
Acknowledgments — This work was supported in part by the Russian Science Foundation (grant No. 22-12-00054). The research was supported in part in the framework of scientific program of the Russian National Center for Physics and Mathematics, topic number 6 “Nuclear and radiation physics” (2023-2025 stage). The authors are grateful to Prof. Yu.Ts. Oganessian for the long-term support of this research.